### v1

# Methods to achieve near-millisecond energy relaxation and dephasing times for a superconducting transmon qubit

Abstract
Superconducting qubits are one of the most promising physical systems for implementing a quantum computer. However, executing quantum algorithms of practical computational advantage requires further improvements in the fidelities of qubit operations, which are currently limited by the energy relaxation and dephasing times of the qubits. Here, we report our measurement results of a high-coherence transmon qubit with energy relaxation and echo dephasing times surpassing those in the existing literature. We measure a qubit frequency of 2.890 GHz, an energy relaxation time with a median of µs and a maximum of µs, and an echo dephasing time with a median of µs and a maximum of µs. We report details of our design, fabrication process, and measurement setup to facilitate the reproduction and wide adoption of high-coherence transmon qubits in the academia and industry.
NOTE ADDED TO arXiv VERSION: We appreciate any early feedback on the manuscript, which we will take into account before the journal submission planned for mid-August 2024.
I Introduction
In the past two decades, the rapid development of superconducting qubits has rendered them one of the most promising candidates for realizing large-scale quantum computers [1, 2, 3, 4, 5]. The transmon qubit, proposed in 2007, has become the most widely used superconducting qubit due to its simplicity and performance [6, 7]. It consists of a Josephson junction and a shunt capacitor which reduces the sensitivity of the qubit to charge noise but keeps the qubit sufficiently anharmonic for realizing fast, high-fidelity gate operations.
Short coherence times have historically been a disadvantage of superconducting qubits, but the previous decade has shown continuous improvements on this challenge, leading to increased fidelities of qubit operations. The longest reported energy relaxation times of transmons have approached but not surpassed 500-µs median value as shown in Table 1 based on Refs. [8, 9, 10, 11, 12, 13, 14]. Recently, there have also been significant improvements in transmon echo dephasing times with the longest reported values extending to 300–400 µs [12, 9, 14]. With other types of superconducting qubits such as 0- qubits and fluxonium qubits, energy relaxation times above a millisecond have been observed, but these qubits have either had short dephasing times or low qubit frequencies [15, 16].
| Ref. | (GHz) | (µs) | (106) | (µs) |
| 8 | 3.73 | 303 | 7.1 | 201 |
| 9 | 3.92 | 401 | 9.9 | N/A |
| 10 | 3.86 | 393 | 9.5 | N/A |
| 11 | 4.24 | 292 | 7.7 | N/A |
| 12 | 3.02 | 270 | 5.1 | 307 |
| 13 | 3.95 | 323 | 8.0 | 286 |
| 14 | 4.79 | 210 | 6.5 | 290 |
| This work | 2.89 | 502 | 9.1 | 541 |
Here, we report our latest results on high-coherence transmon qubits with improved energy relaxation and echo dephasing times. Importantly, we describe in detail our design, fabrication process, and measurement setup with the intention to promote the reproduction of our result by other academic and industrial groups.
II Sample and its fabrication
Figure 1 shows a microscope image of a sample which is identical to the one measured in this work. It contains four transmon qubits –, each of which couples to a coplanar waveguide resonator for readout. Qubits Q1 and Q3 are flux-tunable by the use of a superconducting quantum interference device (SQUID), whereas Q2 and Q4 are fixed-frequency qubits. The four qubit–resonator pairs are identical except for the size of the Josephson junction and the length of the resonator. The readout resonators couple to a shared Purcell filter, which reduces the energy relaxation of the qubits into the readout lines [17, 18]. The layout design file of the sample is available in Ref. [19].
This section provides a detailed description of our sample fabrication process, which is an adaptation of the recipe described in Ref. [14] to the chemicals and equipment available to us. Notable differences include the choice of CF4 as the processing gas for Nb etching, which improves the reproducibility of the etch, and the pre-dicing of the sample between the Al evaporation and lift-off, which minimizes the exposure of the sample to the ambient atmosphere. The pieces of equipment used in our process are listed in Table 2.
II.0.1 Substrate and niobium patterning
The sample is fabricated using a 675-µm-thick 6-inch (100)-oriented high-resistivity (10 kcm) intrinsic-silicon wafer sourced from Siegert Wafer. The pre-cleaning of the wafer follows Ref. [20] and begins with an RCA solvent clean. The wafer then undergoes a dip in dilute hydrofluoric (HF) acid (1:50) for 1 min. After being rapidly transferred to the sputtering tool to minimize exposure to the ambient atmosphere, the wafer is baked at 300 ∘C under vacuum. The sputtering of a 200-nm Nb film is carried out near room temperature and at 2600-W power. The sputter target has a purity of 99.998%. After the sputtering, the wafer is coated with a protective layer of AZ 5214E photoresist and diced into 25 mm 30 mm rectangular coupons with a dicing saw.
The resonators, coplanar waveguides, ground plane, and transmon capacitors are patterned as described below. The selected coupon is sonicated in acetone and isopropyl alcohol (IPA) for 3 min each to remove the protective resist layer and dried with a nitrogen gun. The sample is then dehydrated on a hotplate at 110 ∘C for 1 min, spin coated with AZ 5214E photoresist at 4000 rpm, and baked at 110 ∘C for 1 min to achieve a coating thickness of 1.4 µm. The photoresist is exposed using a maskless aligner with a laser wavelength of 405 nm and a dose of 130 mJcm2. Subsequently, the photoresist is developed in AZ 726MIF for 1.5 min and rinsed in deionized water (DIW) for 1.5 min.
The Nb film is patterned in a plasma processing system by a chemical dry-etching process, clearing the areas revealed by the resist development step. Immediately before the process, the empty plasma chamber is cleaned with a combination of CF4 and O2 gases for a total of 5 min. We then load the sample into the chamber and apply oxygen plasma ashing for 10 s at 100 mTorr chamber pressure, 40 sccm gas flow, and 150 W rf source power to remove the post-development residue of photoresist on the sample surface. Then, we pump out the oxygen and introduce CF4 at 50-mTorr chamber pressure, 20-sccm flow, and 30-W source power. To ensure that the Nb film is fully etched, we carry out an additional etch after unloading, visually inspecting, and re-loading the sample. The etch rate of Si is significantly lower than that of Nb, which helps us to achieve a convenient control of the etch depth. After the etching process, we remove some of the residual chemicals by applying another oxygen plasma ashing for 2 min without breaking the vacuum.
The photoresist is removed by immersing the sample in an -methylpyrrolidone (NMP)-based solvent Remover PG at 80 ∘C overnight and sonicating it in the same Remover PG, acetone, and IPA for 3 min each. Subsequently, the sample is dried with a nitrogen gun and plasma-ashed again for 2 min. Using a profilometer, we measure the etch depth to be 250 nm, which implies that the Si substrate is etched by 50 nm.
II.0.2 Electron-beam lithography
In order to remove the oxide layers on the Nb and Si surfaces, we immerse the sample in dilute HF acid (0.5%) for 10 min and rinse it in DIW for 5 min [21]. The sample is then carried to a spin coater while immersed in fresh DIW to minimize its exposure to the ambient atmosphere. After spin-drying the sample, we immediately spin coat the methyl methacrylate (MMA) EL11 copolymer resist (11% solid content in ethyl lactate) at 4000 rpm, bake it at 180 ∘C for 5 min, and cool it for 3 min. Then we coat the sample with polymethyl methacrylate (PMMA) 950 A4 resist (950,000 molecular weight, 4% solid content in anisole) at 1000 rpm and bake it at 180 ∘C for 5 min. This creates a two-layer resist stack with approximately 500 nm of MMA and 400 nm of PMMA.
The resist mask for Manhattan-style Josephson junctions is patterned onto the resist stack using an electron-beam writer with 100-kV acceleration voltage, 300-µm aperture, and 0.5-nA beam current. We use a dose of 1000 µC/cm2 to define the junction structure in both the PMMA layer and the MMA layer. In addition, we use a dose of 400 µC/cm2 to define an undercut at each end of the line-like structures for the junctions. The undercut serves to separate the aluminum junction evaporated onto the Si substrate from the MMA side walls. We use a mask pattern very similar to that of Ref. [14] to define the Josephson junctions. The electron-beam resist is developed by immersing the sample in methyl isobutyl ketone (MIBK):IPA (1:3) solvent for 5 min, rinsing it in IPA for 1 min, and drying it with a nitrogen gun.
II.0.3 Junction deposition
The Josephson junctions are deposited using an ultra-high-vacuum electron-beam evaporator with separate load-lock, oxidation, and evaporation chambers. After loading the sample, we pump the system for 14 h to reach a load-lock chamber pressure below mBar. Subsequently, we carry out ozone ashing at 10 mBar for 1 min to remove a thin layer of resist residues.
After the ozone cleaning, a high vacuum (10-7 mBar) is pumped again to the load-lock chamber, and the sample is transferred to the oxidation/evaporation chamber. To fabricate the junctions, we evaporate Al at the rate of 0.2 nm/s at tilt and in-plane rotation angles specific for each of the four Al line strips as discussed below. Prior to each Al evaporation step, the oxidation and evaporation chambers are getter-pumped by evaporating Ti with a closed shutter at the rate of 0.1 nm/s for 2 min and waiting for the chamber pressure to decrease below mBar.
For the bottom layer of the junctions, we deposit 40 nm of Al at tilt and planetary angle. Then we create the insulating aluminum oxide layer by static oxidation at 1.2 mbar for 5 min. After the oxidation, we deposit the second Al layer in two steps with tilt and two planetary angles: first 30 nm at and then 30 nm at . This ensures that both sides of the oxidized bottom strip are covered by the second Al layer.
In order to ensure a galvanic contact between the Josephson junctions and the Nb capacitor pads, we transfer the sample to the load-lock chamber and carry out argon milling with tilt at two planetary angles for 2 min each at 10-sccm gas flow, 400-V beam voltage, 60-mA beam current, and 80-V acceleration voltage. This removes any Nb surface oxide that has grown in the areas exposed after the electron-beam resist development [22, 23]. After the argon milling, the sample is transferred back to the oxidation/evaporation chamber, where we deposit the connecting leads between the junction and the Nb pads in three steps with 30/60/60-nm thicknesses, tilt and planetary angles.
As the final step, the sample is transferred back to the load-lock chamber and oxidized at 20 mBar for 10 min to create a clean oxide layer on top of the junctions before exposing the sample to the ambient atmosphere.
II.0.4 Dicing and liftoff
After the Josephson junction deposition, we coat the sample with a protective layer of AZ 5214E photoresist and pre-dice the sample with a dicing saw, cutting two-thirds deep of the substrate thickness. Before attaching the sample to the dicing tape, an ionizer fan is applied to the tape to reduce the possibility of electrostatic-discharge damage on the sample. By pre-dicing the sample prior to the liftoff, we can skip one of the resist removal steps carried out in Ref. [14].
The pre-diced sample is immersed in Remover PG at 80 ∘C for 3 h, after which large Al flakes can be removed from the beaker using a pipette. The sample is sonicated in the same Remover PG, acetone, and IPA for 3 min each. From IPA, the sample is quickly dried with a nitrogen gun and immediately transferred to a vapor prime oven that applies a monolayer of hexamethyldisilazane (HMDS) on the sample surface, which may help to slow down the post-fabrication oxidation of the Nb and Si surfaces. The oven also has the effect of annealing the sample at 150∘C for a total of 20 min under vacuum and nitrogen environments.
| Equipment | Description |
| Sputtering system | Eclipse Mark IV |
| Maskless aligner | Heidelberg Instruments MLA150 |
| Plasma processing system | Oxford Instruments Plasmalab |
| 80 Plus | |
| Electron-beam writer | Raith EBPG5200 |
| Electron-beam evaporator | Plassys MEB700S2-III |
| Vapor prime oven | YES III |
The critical currents of the fabricated junctions are estimated by measuring their room-temperature resistances using a probe station. The critical currents for the sample measured in this work are significantly smaller than we had targeted, which led to low ratios, as low as 20 for qubit Q2. The sample is then manually cleaved into separate chips utilizing the cuts established with the dicing saw, and selected chips are taken for further characterization.
III Setup for qubit measurements
Our experimental setup used for qubit characterization and measurements is presented in Fig. 2, with detailed information on measurement equipment and components provided in Table 3. The sample is wire-bonded to a QCage.24 sample holder and placed inside QCage Magnetic Shielding, both of which are supplied by QDevil (under Quantum Machines). The sample is exposed to ambient atmosphere at room temperature for a total of 7 days. It is then cooled down to approximately 10 mK in a Bluefors dilution refrigerator equipped with a tin-plated copper shield at 10 mK and an Amumetal magnetic shield just inside the outer vacuum chamber (OVC). We directly generate the qubit control and readout signals without any analog mixers by using a Xilinx RFSoC evaluation board with QICK firmware [24, 25] locked to a Rb frequency standard. The signals then pass through rf attenuators, an Eccosorb infrared filter, and a low pass filter before reaching the sample. The readout signal coming out of the sample goes through a three-wave-mixing Josephson traveling-wave parametric amplifier (TWPA), a cryogenic high-electron-mobility-transistor (HEMT) amplifier, and a room-temperature HEMT amplifier before being digitized by RFSoC. During the first cooldown, the TWPA is not pumped and therefore does not provide any amplification.
| Equipment | Description | Abbreviation |
| Dilution refrigerator | Bluefors XLD400sl | |
| RFSoC evaluation board | Xilinx Zynq UltraScale+ RFSoC ZCU216 with add-ons CLK104 and XM655 | |
| Frequency standard | Stanford Research Systems FS725 10-MHz Rb Standard | |
| Coaxial cables | Totoku TCF258AA2000/1500 (room temperature) | |
| Mini-Circuits 141-72SM+ (room temperature) | ||
| Bluefors 0.86 mm SCuNi-CuNi semi-rigid (cryogenic input lines) | ||
| Bluefors 0.86 mm NbTi-NbTi semi-rigid (cryogenic output lines, MXC to 4 K) | ||
| Bluefors 2.19 mm SCuNi-CuNi semi-rigid (cryogenic output lines, 4 K to RT) | ||
| HEMT amplifiers | Low Noise Factory LNF-LNR4_8F (room temperature) | HEMT |
| Low Noise Factory LNF-LNC4_8F (cryogenic) | HEMT∗ | |
| Low pass filters | RLC F-30-8000-R (3-dB point = 8.4 GHz, 60-dB stopband = 10.8–40 GHz) | LPF |
| Mini-Circuits VLF-5850+ (3-dB point = 6.54 GHz) | LPF∗ | |
[…]
| Bandpass filters | Mini-Circuits VBFZ-3590+ (2-dB passband = 3–4.3 GHz) | VBFZ |
| Mini-Circuits VBFZ-6260-S+ (2-dB passband = 5.6–7 GHz) | VBFZ∗ | |
| Highpass filter | Mini-Circuits VHF-8400+ (20-dB stopband = DC–7.8 GHz) | HPF |
| Eccosorb filter | Kawashima Manufacturing 3.5mm Filter Connector L9.6 (0.32 dB/GHz) | Ecco. |
| Attenuator | Bluefors cryogenic attenuators | dB |
| XMA 2082-6418-20-CRYO | 20 dB∗ | |
| Mini-Circuits BW-S10-2W263+ (room temperature) | 10 dB† | |
| Isolators | Low Noise Factory LNF-ISISC4_8A (4–8 GHz) | |
| Low Noise Factory LNF-ISISC4_12A (4–12 GHz) | ||
| Circulator | Low Noise Factory LNF-CIC4_12A (4–12 GHz) | |
| TWPA | Supplied by VTT (Technical Research Center of Finland Ltd.) | TWPA |
| TWPA pump generator | Valon 5015 Frequency Synthesizer | |
| Voltage source | Stanford Research Systems SIM928 Isolated Voltage Source | V |
| Sample holder | QDevil QCage.24 with QCage Magnetic Shielding | QCage.24 |
After the first cooldown, we reconfigure the attenuators as described in the caption of Fig. 2 before starting another cooldown. During the second cooldown, we compare measurements with and without a pump signal applied to the TWPA. The 10.5-GHz pump signal passes through a 4–12 GHz circulator, reflects from a 8-GHz lowpass filter, and travels through the circulator again before reaching the TWPA. This configuration allows us to combine the pump and readout signals without the power dissipation of a directional coupler or the impedance mismatch of a diplexer.
IV Measurement Results
We measure the fundamental modes of the four readout resonators to have frequencies ranging from GHz to GHz and the fixed-frequency qubits to have qubit frequencies of GHz and GHz. The flux-tunable qubits have maximum frequencies of GHz and GHz which we find by sweeping the external magnetic flux applied to the sample. We then follow the standard procedure for measuring the energy relaxation time and echo dephasing time of each qubit [2].
| Qubit | (GHz) | (µs) | (µs) | (106) |
| (tunable) | 4.047 | 195 | 237 | 5.0 |
| (fixed) | 2.890 | 502 | 549 | 9.1 |
| (tunable) | 4.593 | 44 | 28 | 1.3 |
| (fixed) | 3.295 | 154 | 114 | 3.2 |
During the first cooldown, we measure that the fixed-frequency qubit exhibits especially long energy relaxation and echo dephasing times: medians of µs and of µs and maxima of µs and µs. The data for these results are shown in Fig. 3 in the form of distributions for repeated measurements of and and time traces for the maximum values. In each time trace, we repeat and average 500 measurements for each value of delay time between a -pulse and a readout pulse. The number of repetitions is for the measurement. The measurement results for all four qubits during the first cooldown is summarized in Table 4.
During the second cooldown, we carry out a more extensive measurement of the qubit . The qubit frequency is measured to have decreased by MHz compared to the first cooldown. The readout resonator of this qubit is measured to have a linewidth of 2.63 MHz and a dispersive shift of MHz. Then, we measure the energy relaxation and echo dephasing times with and without the pump signal applied to the TWPA. We fit an exponential function to the obtained data points, each of which is an average of 500 measurements, and compute the relative uncertainties of the fit parameters. We find that the TWPA does not significantly affect the coherence times but improves the certainties of the and obtained from the fit. These results are presented in Fig. 4. We exclude from further analysis the values with relative uncertainties exceeding and values with relative uncertainties exceeding . The resulting data are shown in Fig. 6. We obtain medians of µs and µs and maxima of µs and µs. Figure 5 shows the time traces for the maximum values.
The data for the coherence time measurements carried out in this section are available in Ref. [19].
V Discussion
In this work, we presented detailed information on our design, fabrication methods, and measurement setup for a high-coherence transmon qubit. One of the four qubits on the measured sample achieved a median energy relaxation time of µs and a median echo dephasing time of µs. These results surpass the previous results for a transmon qubit reported in the literature. However, the and were significantly shorter in the second cooldown. This may be due to the redistribution of environmental fluctuators and oxidation of the sample surface. Our fabrication method and experimental setup can also be applied to other types of superconducting qubits such as the unimon qubit to enhance its energy relaxation and dephasing times [26].
In conclusion, this result represents a significant step in the development of high-coherence superconducting qubits by approaching the millisecond mark for the energy relaxation and dephasing times of a transmon qubit. It also demonstrates the robustness and reproducibility of the fabrication recipe described in Ref. 14 and the effectiveness of the equipment and components listed in Tables 2 and 3. Detailed reporting of a high-coherence qubit will benefit the research community and accelerate the global efforts on developing quantum sensors, quantum simulators, and quantum computers based on superconducting quantum technologies.
Acknowledgements.
We acknowledge the Research Council of Finland under its Centre of Excellence Quantum Technology Finland (Grant nos. 336810, 352925, 352934 and 352935), the European Research Council under the Advanced Grant ConceptQ (Grant no. 101053801), the Finnish Quantum Institute InstituteQ, and the provision of facilities and technical support by Aalto University at the OtaNano-Micronova Nanofabrication Centre.References
- Kjaergaard et al. [2020] M. Kjaergaard, M. E. Schwartz, J. Braumüller, P. Krantz, J. I.-J. Wang, S. Gustavsson, and W. D. Oliver, Superconducting Qubits: Current State of Play, Annual Review of Condensed Matter Physics 11, 369 (2020).
- Krantz et al. [2019] P. Krantz, M. Kjaergaard, F. Yan, T. P. Orlando, S. Gustavsson, and W. D. Oliver, A quantum engineer’s guide to superconducting qubits, Applied Physics Reviews 6, 021318 (2019).
- Mamgain et al. [2023] A. Mamgain, S. S. Khaire, U. Singhal, I. Ahmad, L. A. Patel, K. D. Helambe, S. Majumder, V. Singh, and B. Suri, A Review of Developments in Superconducting Quantum Processors, Journal of the Indian Institute of Science 103, 633 (2023).
- Huang et al. [2020] H.-L. Huang, D. Wu, D. Fan, and X. Zhu, Superconducting Quantum Computing: A Review, Science China Information Sciences 63, 180501 (2020), arXiv:2006.10433 [quant-ph].
- Ezratty [2023] O. Ezratty, Perspective on superconducting qubit quantum computing, The European Physical Journal A 59, 94 (2023).
- Koch et al. [2007] J. Koch, T. M. Yu, J. Gambetta, A. A. Houck, D. I. Schuster, J. Majer, A. Blais, M. H. Devoret, S. M. Girvin, and R. J. Schoelkopf, Charge-insensitive qubit design derived from the Cooper pair box, Physical Review A 76, 042319 (2007).
- Blais et al. [2021] A. Blais, A. L. Grimsmo, S. Girvin, and A. Wallraff, Circuit quantum electrodynamics, Reviews of Modern Physics 93, 025005 (2021).
- Place et al. [2021] A. P. M. Place, L. V. H. Rodgers, P. Mundada, B. M. Smitham, M. Fitzpatrick, Z. Leng, A. Premkumar, J. Bryon, A. Vrajitoarea, S. Sussman, G. Cheng, T. Madhavan, H. K. Babla, X. H. Le, Y. Gang, B. Jäck, A. Gyenis, N. Yao, R. J. Cava, N. P. De Leon, and A. A. Houck, New material platform for superconducting transmon qubits with coherence times exceeding 0.3 milliseconds, Nature Communications 12, 1779 (2021).
- Wang et al. [2022] C. Wang, X. Li, H. Xu, Z. Li, J. Wang, Z. Yang, Z. Mi, X. Liang, T. Su, C. Yang, G. Wang, W. Wang, Y. Li, M. Chen, C. Li, K. Linghu, J. Han, Y. Zhang, Y. Feng, Y. Song, T. Ma, J. Zhang, R. Wang, P. Zhao, W. Liu, G. Xue, Y. Jin, and H. Yu, Towards practical quantum computers: transmon qubit with a lifetime approaching 0.5 milliseconds, npj Quantum Information 8, 3 (2022).
- Gordon et al. [2022] R. T. Gordon, C. E. Murray, C. Kurter, M. Sandberg, S. A. Hall, K. Balakrishnan, R. Shelby, B. Wacaser, A. A. Stabile, J. W. Sleight, M. Brink, M. B. Rothwell, K. P. Rodbell, O. Dial, and M. Steffen, Environmental radiation impact on lifetimes and quasiparticle tunneling rates of fixed-frequency transmon qubits, Applied Physics Letters 120, 074002 (2022).
- Deng et al. [2023] H. Deng, Z. Song, R. Gao, T. Xia, F. Bao, X. Jiang, H.-S. Ku, Z. Li, X. Ma, J. Qin, H. Sun, C. Tang, T. Wang, F. Wu, W. Yu, G. Zhang, X. Zhang, J. Zhou, X. Zhu, Y. Shi, H.-H. Zhao, and C. Deng, Titanium Nitride Film on Sapphire Substrate with Low Dielectric Loss for Superconducting Qubits, Physical Review Applied 19, 024013 (2023).
- Biznárová et al. [2023] J. Biznárová, A. Osman, E. Rehnman, L. Chayanun, C. Križan, P. Malmberg, M. Rommel, C. Warren, P. Delsing, A. Yurgens, J. Bylander, and A. F. Roudsari, Mitigation of interfacial dielectric loss in aluminum-on-silicon superconducting qubits (2023), arXiv:2310.06797 [cond-mat, physics:quant-ph].
- Bal et al. [2024] M. Bal, A. A. Murthy, S. Zhu, F. Crisa, X. You, Z. Huang, T. Roy, J. Lee, D. V. Zanten, R. Pilipenko, I. Nekrashevich, A. Lunin, D. Bafia, Y. Krasnikova, C. J. Kopas, E. O. Lachman, D. Miller, J. Y. Mutus, M. J. Reagor, H. Cansizoglu, J. Marshall, D. P. Pappas, K. Vu, K. Yadavalli, J.-S. Oh, L. Zhou, M. J. Kramer, F. Lecocq, D. P. Goronzy, C. G. Torres-Castanedo, P. G. Pritchard, V. P. Dravid, J. M. Rondinelli, M. J. Bedzyk, M. C. Hersam, J. Zasadzinski, J. Koch, J. A. Sauls, A. Romanenko, and A. Grassellino, Systematic improvements in transmon qubit coherence enabled by niobium surface encapsulation, npj Quantum Information 10, 43 (2024).
- Kono et al. [2024] S. Kono, J. Pan, M. Chegnizadeh, X. Wang, A. Youssefi, M. Scigliuzzo, and T. J. Kippenberg, Mechanically induced correlated errors on superconducting qubits with relaxation times exceeding 0.4 ms, Nature Communications 15, 3950 (2024).
- Gyenis et al. [2021] A. Gyenis, P. S. Mundada, A. Di Paolo, T. M. Hazard, X. You, D. I. Schuster, J. Koch, A. Blais, and A. A. Houck, Experimental realization of an intrinsically error-protected superconducting qubit, PRX Quantum 2, 010339 (2021), arXiv:1910.07542 [cond-mat, physics:quant-ph].
- Somoroff et al. [2023] A. Somoroff, Q. Ficheux, R. A. Mencia, H. Xiong, R. Kuzmin, and V. E. Manucharyan, Millisecond Coherence in a Superconducting Qubit, Physical Review Letters 130, 267001 (2023).
- Jeffrey et al. [2014] E. Jeffrey, D. Sank, J. Mutus, T. White, J. Kelly, R. Barends, Y. Chen, Z. Chen, B. Chiaro, A. Dunsworth, A. Megrant, P. O’Malley, C. Neill, P. Roushan, A. Vainsencher, J. Wenner, A. Cleland, and J. M. Martinis, Fast Accurate State Measurement with Superconducting Qubits, Physical Review Letters 112, 190504 (2014).
- Sete et al. [2015] E. A. Sete, J. M. Martinis, and A. N. Korotkov, Quantum theory of a bandpass Purcell filter for qubit readout, Physical Review A 92, 012325 (2015).
- Tuokkola et al. [2024] M. Tuokkola, Y. Sunada, H. Kivijärvi, L. Grönberg, J.-P. Kaikkonen, V. Vesterinen, J. Govenius, and M. Möttönen, Data and GDS file for ”Methods to achieve near-millisecond energy relaxation and dephasing times for a superconducting transmon qubit (2024).
- Burnett et al. [2018] J. Burnett, A. Bengtsson, D. Niepce, and J. Bylander, Noise and loss of superconducting aluminium resonators at single photon energies, Journal of Physics: Conference Series 969, 012131 (2018).
- Verjauw et al. [2021] J. Verjauw, A. Potočnik, M. Mongillo, R. Acharya, F. Mohiyaddin, G. Simion, A. Pacco, T. Ivanov, D. Wan, A. Vanleenhove, L. Souriau, J. Jussot, A. Thiam, J. Swerts, X. Piao, S. Couet, M. Heyns, B. Govoreanu, and I. Radu, Investigation of Microwave Loss Induced by Oxide Regrowth in High- Q Niobium Resonators, Physical Review Applied 16, 014018 (2021).
- Osman et al. [2021] A. Osman, J. Simon, A. Bengtsson, S. Kosen, P. Krantz, D. P. Lozano, M. Scigliuzzo, P. Delsing, J. Bylander, and A. Fadavi Roudsari, Simplified Josephson-junction fabrication process for reproducibly high-performance superconducting qubits, Applied Physics Letters 118, 064002 (2021).
- Bilmes et al. [2021] A. Bilmes, A. K. Händel, S. Volosheniuk, A. V. Ustinov, and J. Lisenfeld, In-situ bandaged Josephson junctions for superconducting quantum processors, Superconductor Science and Technology 34, 125011 (2021).
- Stefanazzi et al. [2022] L. Stefanazzi, K. Treptow, N. Wilcer, C. Stoughton, C. Bradford, S. Uemura, S. Zorzetti, S. Montella, G. Cancelo, S. Sussman, A. Houck, S. Saxena, H. Arnaldi, A. Agrawal, H. Zhang, C. Ding, and D. I. Schuster, The QICK (Quantum Instrumentation Control Kit): Readout and control for qubits and detectors, Review of Scientific Instruments 93, 044709 (2022).
- Ding et al. [2024] C. Ding, M. Di Federico, M. Hatridge, A. Houck, S. Leger, J. Martinez, C. Miao, D. S. I, L. Stefanazzi, C. Stoughton, S. Sussman, K. Treptow, S. Uemura, N. Wilcer, H. Zhang, C. Zhou, and G. Cancelo, Experimental advances with the QICK (Quantum Instrumentation Control Kit) for superconducting quantum hardware, Physical Review Research 6, 013305 (2024).
- Hyyppä et al. [2022] E. Hyyppä, S. Kundu, C. F. Chan, A. Gunyhó, J. Hotari, D. Janzso, K. Juliusson, O. Kiuru, J. Kotilahti, A. Landra, W. Liu, F. Marxer, A. Mäkinen, J.-L. Orgiazzi, M. Palma, M. Savytskyi, F. Tosto, J. Tuorila, V. Vadimov, T. Li, C. Ockeloen-Korppi, J. Heinsoo, K. Y. Tan, J. Hassel, and M. Möttönen, Unimon qubit, Nature Communications 13, 6895 (2022).
Appendix A Extended Data Figures

### v2

@@
+Methods to achieve near-millisecond energy relaxation and dephasing times for a superconducting transmon qubit Tiivistelmä
 Methods to achieve near-millisecond energy relaxation and dephasing times for a superconducting transmon qubit
-Abstract
-Superconducting qubits are one of the most promising physical systems for implementing a quantum computer.
+Tiivistelmä
+Superconducting qubits are one of the most promising physical systems for implementing quantum computers.
 However, executing quantum algorithms of practical computational advantage requires further improvements in the fidelities of qubit operations, which are currently limited by the energy relaxation and dephasing times of the qubits.
@@
 We measure a qubit frequency of 2.890 GHz, an energy relaxation time with a median of µs and a maximum of µs, and an echo dephasing time with a median of µs and a maximum of µs.
-We report details of our design, fabrication process, and measurement setup to facilitate the reproduction and wide adoption of high-coherence transmon qubits in the academia and industry.
-NOTE ADDED TO arXiv VERSION: We appreciate any early feedback on the manuscript, which we will take into account before the journal submission planned for mid-August 2024.
+We report in detail our design, fabrication process, and measurement setup to facilitate the reproduction and wide adoption of high-coherence transmon qubits in the academia and industry.
 I Introduction
-In the past two decades, the rapid development of superconducting qubits has rendered them one of the most promising candidates for realizing large-scale quantum computers [1, 2, 3, 4, 5].
+In the past two decades, the rapid development of superconducting qubits has rendered them one of the most promising candidates for realizing large-scale quantum computers [1, 2, 3, 4, 5, 6].
 The transmon qubit, proposed in 2007, has become the most widely used superconducting qubit due to its simplicity and performance [7, 8].
@@
 Short coherence times have historically been a disadvantage of superconducting qubits, but the previous decade has shown continuous improvements on this challenge, leading to increased fidelities of qubit operations.
-The longest reported energy relaxation times of transmons have approached but not surpassed 500-µs median value as shown in Table 1 based on Refs.
-[8, 9, 10, 11, 12, 13, 14].
-Recently, there have also been significant improvements in transmon echo dephasing times with the longest reported values extending to 300–400 µs [12, 9, 14].
-With other types of superconducting qubits such as 0- qubits and fluxonium qubits, energy relaxation times above a millisecond have been observed, but these qubits have either had short dephasing times or low qubit frequencies [15, 16].
-| Ref. | (GHz) | (µs) | (106) | (µs) |
-| 8 | 3.73 | 303 | 7.1 | 201 |
-| 9 | 3.92 | 401 | 9.9 | N/A |
-| 10 | 3.86 | 393 | 9.5 | N/A |
-| 11 | 4.24 | 292 | 7.7 | N/A |
-| 12 | 3.02 | 270 | 5.1 | 307 |
-| 13 | 3.95 | 323 | 8.0 | 286 |
-| 14 | 4.79 | 210 | 6.5 | 290 |
-| This work | 2.89 | 502 | 9.1 | 541 |
+The longest reported energy relaxation times of transmons have approached but not surpassed 500-µs median value as shown in Fig.
+1 and Extended Data Table 1 based on Refs.
+[9, 10, 11, 12, 13, 14, 15, 16].
+Recently, there have also been significant improvements in transmon echo dephasing times with the longest reported average extending to 307 µs [10, 13, 14, 16].
+Transmon qubits with coherence times above 100 µs have also been reported, for example, in Refs.
+[17, 18, 19, 20, 21, 22].
+With other types of superconducting qubits such as 0- qubits and fluxonium qubits, energy relaxation times above a millisecond have been observed, but these qubits have either had short dephasing times or low qubit frequencies [23, 24, 25].
 Here, we report our latest results on high-coherence transmon qubits with improved energy relaxation and echo dephasing times.
@@
 [28].
+For more details of the sample, see Extended Data Figs.
+1 and 2
 This section provides a detailed description of our sample fabrication process, which is an adaptation of the recipe described in Ref.
@@
 Notable differences include the choice of CF4 as the processing gas for Nb etching, which improves the reproducibility of the etch, and the pre-dicing of the sample between the Al evaporation and lift-off, which minimizes the exposure of the sample to the ambient atmosphere.
-The pieces of equipment used in our process are listed in Table 2.
+The pieces of equipment used in our process are listed in Extended Data Table 2.
 II.0.1 Substrate and niobium patterning
@@
 [29] and begins with an RCA solvent clean.
-The wafer then undergoes a dip in dilute hydrofluoric (HF) acid (1:50) for 1 min.
+The wafer then undergoes a dip in dilute hydrofluoric (HF) acid (1:100) for 1 min.
 After being rapidly transferred to the sputtering tool to minimize exposure to the ambient atmosphere, the wafer is baked at 300 ∘C under vacuum.
@@
 The oven also has the effect of annealing the sample at 150∘C for a total of 20 min under vacuum and nitrogen environments.
-| Equipment | Description |
-| Sputtering system | Eclipse Mark IV |
-| Maskless aligner | Heidelberg Instruments MLA150 |
-| Plasma processing system | Oxford Instruments Plasmalab |
-| 80 Plus | |
-| Electron-beam writer | Raith EBPG5200 |
-| Electron-beam evaporator | Plassys MEB700S2-III |
-| Vapor prime oven | YES III |
 The critical currents of the fabricated junctions are estimated by measuring their room-temperature resistances using a probe station.
@@
 Our experimental setup used for qubit characterization and measurements is presented in Fig.
-2, with detailed information on measurement equipment and components provided in Table 3.
+3, with detailed information on measurement equipment and components provided in Extended Data Table 3.
 The sample is wire-bonded to a QCage.24 sample holder and placed inside QCage Magnetic Shielding, both of which are supplied by QDevil (under Quantum Machines).
@@
 During the first cooldown, the TWPA is not pumped and therefore does not provide any amplification.
-| Equipment | Description | Abbreviation |
-| Dilution refrigerator | Bluefors XLD400sl | |
-| RFSoC evaluation board | Xilinx Zynq UltraScale+ RFSoC ZCU216 with add-ons CLK104 and XM655 | |
-| Frequency standard | Stanford Research Systems FS725 10-MHz Rb Standard | |
-| Coaxial cables | Totoku TCF258AA2000/1500 (room temperature) | |
-| Mini-Circuits 141-72SM+ (room temperature) | ||
-| Bluefors 0.86 mm SCuNi-CuNi semi-rigid (cryogenic input lines) | ||
-| Bluefors 0.86 mm NbTi-NbTi semi-rigid (cryogenic output lines, MXC to 4 K) | ||
-| Bluefors 2.19 mm SCuNi-CuNi semi-rigid (cryogenic output lines, 4 K to RT) | ||
-| HEMT amplifiers | Low Noise Factory LNF-LNR4_8F (room temperature) | HEMT |
-| Low Noise Factory LNF-LNC4_8F (cryogenic) | HEMT∗ | |
-| Low pass filters | RLC F-30-8000-R (3-dB point = 8.4 GHz, 60-dB stopband = 10.8–40 GHz) | LPF |
-| Mini-Circuits VLF-5850+ (3-dB point = 6.54 GHz) | LPF∗ | |
-[…]
-| Bandpass filters | Mini-Circuits VBFZ-3590+ (2-dB passband = 3–4.3 GHz) | VBFZ |
-| Mini-Circuits VBFZ-6260-S+ (2-dB passband = 5.6–7 GHz) | VBFZ∗ | |
-| Highpass filter | Mini-Circuits VHF-8400+ (20-dB stopband = DC–7.8 GHz) | HPF |
-| Eccosorb filter | Kawashima Manufacturing 3.5mm Filter Connector L9.6 (0.32 dB/GHz) | Ecco. |
-| Attenuator | Bluefors cryogenic attenuators | dB |
-| XMA 2082-6418-20-CRYO | 20 dB∗ | |
-| Mini-Circuits BW-S10-2W263+ (room temperature) | 10 dB† | |
-| Isolators | Low Noise Factory LNF-ISISC4_8A (4–8 GHz) | |
-| Low Noise Factory LNF-ISISC4_12A (4–12 GHz) | ||
-| Circulator | Low Noise Factory LNF-CIC4_12A (4–12 GHz) | |
-| TWPA | Supplied by VTT (Technical Research Center of Finland Ltd.) | TWPA |
-| TWPA pump generator | Valon 5015 Frequency Synthesizer | |
-| Voltage source | Stanford Research Systems SIM928 Isolated Voltage Source | V |
-| Sample holder | QDevil QCage.24 with QCage Magnetic Shielding | QCage.24 |
+| Qubit | (GHz) | (µs) | (µs) | (106) |
+| (tunable) | 4.047 | 195 | 237 | 5.0 |
+| (fixed) | 2.890 | 502 | 549 | 9.1 |
+| (tunable) | 4.593 | 44 | 28 | 1.3 |
+| (fixed) | 3.295 | 154 | 114 | 3.2 |
 After the first cooldown, we reconfigure the attenuators as described in the caption of Fig.
@@
 We then follow the standard procedure for measuring the energy relaxation time and echo dephasing time of each qubit [2].
-| Qubit | (GHz) | (µs) | (µs) | (106) |
-| (tunable) | 4.047 | 195 | 237 | 5.0 |
-| (fixed) | 2.890 | 502 | 549 | 9.1 |
-| (tunable) | 4.593 | 44 | 28 | 1.3 |
-| (fixed) | 3.295 | 154 | 114 | 3.2 |
 During the first cooldown, we measure that the fixed-frequency qubit exhibits especially long energy relaxation and echo dephasing times: medians of µs and of µs and maxima of µs and µs.
 The data for these results are shown in Fig.
-3 in the form of distributions for repeated measurements of and and time traces for the maximum values.
+4(a, b, e, f) in the form of distributions for repeated measurements of and and time traces for the maximum values.
 In each time trace, we repeat and average 500 measurements for each value of delay time between a -pulse and a readout pulse.
@@
 We find that the TWPA does not significantly affect the coherence times but improves the certainties of the and obtained from the fit.
-These results are presented in Fig.
-4.
+These results are presented in Extended Data Fig.
+3.
 We exclude from further analysis the values with relative uncertainties exceeding and values with relative uncertainties exceeding .
 The resulting data are shown in Fig.
-6.
+4(c, d, g, h) and Extended Data Fig.
+4.
 We obtain medians of µs and µs and maxima of µs and µs.
-Figure 5 shows the time traces for the maximum values.
+The time trace for the longest measured deviates from an exponential function, likely because of the qubit frequency shifting during the sweep owing to the charge noise.
 The data for the coherence time measurements carried out in this section are available in Ref. [28].
@@
 It also demonstrates the robustness and reproducibility of the fabrication recipe described in Ref.
-14 and the effectiveness of the equipment and components listed in Tables 2 and 3.
+16 and the effectiveness of the equipment and components listed in Extended Data Tables 2 and 3.
 Detailed reporting of a high-coherence qubit will benefit the research community and accelerate the global efforts on developing quantum sensors, quantum simulators, and quantum computers based on superconducting quantum technologies.
+Data availability
+The data that support the findings of this study are available at https://doi.org/10.5281/zenodo.12819934.
 Acknowledgements.
 We acknowledge the Research Council of Finland under its Centre of Excellence Quantum Technology Finland (Grant nos.
-336810, 352925, 352934 and 352935), the European Research Council under the Advanced Grant ConceptQ (Grant no.
-101053801), the Finnish Quantum Institute InstituteQ, and the provision of facilities and technical support by Aalto University at the OtaNano-Micronova Nanofabrication Centre.References
+336810, 352925, 352934 and 352935) and the Finnish Quantum Flagship, the European Research Council under the Advanced Grant ConceptQ (no.
+101053801), the Finnish Quantum Institute InstituteQ, Business Finland through Quantum Technologies Industrial project (no.
+41419/31/2020), Horizon Europe programme HORIZON-CL4-2022-QUANTUM-01-SGA through OpenSuperQPlus100 project (no.
+101113946), Jane and Aatos Erkko Foundation, QDOC Doctoral Pilot Programme, and the provision of facilities and technical support by Aalto University at the OtaNano-Micronova Nanofabrication Centre.Competing interests
+M.M. declares that he is a Co-Founder and Shareholder of the quantum-computer company IQM Finland and of the quantum-algorithm company QMill Oy. Other authors declare no competing interests.
+Viitteet
 - Kjaergaard et al.
@@
 Suri, A Review of Developments in Superconducting Quantum Processors, Journal of the Indian Institute of Science 103, 633 (2023).
-- Huang et al. [2020] H.-L. Huang, D. Wu, D. Fan, and X. Zhu, Superconducting Quantum Computing: A Review, Science China Information Sciences 63, 180501 (2020), arXiv:2006.10433 [quant-ph].
+- Huang et al. [2020] H.-L. Huang, D. Wu, D. Fan, and X. Zhu, Superconducting quantum computing: a review, Science China Information Sciences 63, 180501 (2020).
 - Ezratty [2023] O. Ezratty, Perspective on superconducting qubit quantum computing, The European Physical Journal A 59, 94 (2023).
+- Acharya et al.
+[2024] R.
+Acharya, L.
+Aghababaie-Beni, I.
+Aleiner, T.
+I.
+Andersen, M.
+Ansmann, F.
+Arute, K.
+Arya, A.
+Asfaw, N.
+Astrakhantsev, J.
+Atalaya, R.
+Babbush, D.
+Bacon, B.
+Ballard, J.
+C.
+Bardin, J.
+Bausch, A.
+Bengtsson, A.
+Bilmes, S.
+Blackwell, S.
+Boixo, G.
+Bortoli, A.
+Bourassa, J.
+Bovaird, L.
+Brill, M.
+Broughton, D.
+A.
+Browne, B.
+Buchea, B.
+B.
+Buckley, D.
+A.
+Buell, T.
+Burger, B.
+Burkett, N.
+Bushnell, A.
+Cabrera, J.
+Campero, H.-S.
+Chang, Y.
+Chen, Z.
+Chen, B.
+Chiaro, D.
+Chik, C.
+Chou, J.
+Claes, A.
+Y.
+Cleland, J.
+Cogan, R.
+Collins, P.
+Conner, W.
+Courtney, A.
+L.
+Crook, B.
+Curtin, S.
+Das, A.
+Davies, L.
+De Lorenzo, D.
+M.
+Debroy, S.
+Demura, M.
+Devoret, A.
+Di Paolo, P.
+Donohoe, I.
+Drozdov, A.
+Dunsworth, C.
+Earle, T.
+Edlich, A.
+Eickbusch, A.
+M.
+Elbag, M.
+Elzouka, C.
+Erickson, L.
+Faoro, E.
+Farhi, V.
+S.
+Ferreira, L.
+F.
+Burgos, E.
+Forati, A.
+G.
+Fowler, B.
+Foxen, S.
+Ganjam, G.
+Garcia, R.
+Gasca, E.
+Genois, W.
+Giang, C.
+Gidney, D.
+Gilboa, R.
+Gosula, A.
+G.
+Dau, D.
+Graumann, A.
+Greene, J.
+A.
+Gross, S.
+Habegger, J.
+Hall, M.
+C.
+Hamilton, M.
+Hansen, M.
+P.
+Harrigan, S.
+D.
+Harrington, F.
+J.
+H.
+Heras, S.
+Heslin, P.
+Heu, O.
+Higgott, G.
+Hill, J.
+Hilton, G.
+Holland, S.
+Hong, H.-Y.
+Huang, A.
+Huff, W.
+J.
+Huggins, L.
+B.
+Ioffe, S.
+V.
+Isakov, J.
+Iveland, E.
+Jeffrey, Z.
+Jiang, C.
+Jones, S.
+Jordan, C.
+Joshi, P.
+Juhas, D.
+Kafri, H.
+Kang, A.
+H.
+Karamlou, K.
+Kechedzhi, J.
+Kelly, T.
+Khaire, T.
+Khattar, M.
+Khezri, S.
+Kim, P.
+V.
+Klimov, A.
+R.
+Klots, B.
+Kobrin, P.
+Kohli, A.
+N.
+Korotkov, F.
+Kostritsa, R.
+Kothari, B.
+Kozlovskii, J.
+M.
+Kreikebaum, V.
+D.
+Kurilovich, N.
+Lacroix, D.
+Landhuis, T.
+Lange-Dei, B.
+W.
+Langley, P.
+Laptev, K.-M.
+Lau, L.
+L.
+Guevel, J.
+Ledford, K.
+Lee, Y.
+D.
+Lensky, S.
+Leon, B.
+J.
+Lester, W.
+Y.
+Li, Y.
+Li, A.
+T.
+Lill, W.
+Liu, W.
+P.
+Livingston, A.
+Locharla, E.
+Lucero, D.
+Lundahl, A.
+Lunt, S.
+Madhuk, F.
+D.
+Malone, A.
+Maloney, S.
+Mandrá, L.
+S.
+Martin, S.
+Martin, O.
+Martin, C.
+Maxfield, J.
+R.
+McClean, M.
+McEwen, S.
+Meeks, A.
+Megrant, X.
+Mi, K.
+C.
+Miao, A.
+Mieszala, R.
+Molavi, S.
+Molina, S.
+Montazeri, A.
+Morvan, R.
+Movassagh, W.
+Mruczkiewicz, O.
+Naaman, M.
+Neeley, C.
+Neill, A.
+Nersisyan, H.
+Neven, M.
+Newman, J.
+H.
+Ng, A.
+Nguyen, M.
+Nguyen, C.-H.
+Ni, T.
+E.
+O’Brien, W.
+D.
+Oliver, A.
+Opremcak, K.
+Ottosson, A.
+Petukhov, A.
+Pizzuto, J.
+Platt, R.
+Potter, O.
+Pritchard, L.
+P.
+Pryadko, C.
+Quintana, G.
+Ramachandran, M.
+J.
+Reagor, D.
+M.
+Rhodes, G.
+Roberts, E.
+Rosenberg, E.
+Rosenfeld, P.
+Roushan, N.
+C.
+Rubin, N.
+Saei, D.
+Sank, K.
+Sankaragomathi, K.
+J.
+Satzinger, H.
+F.
+Schurkus, C.
+Schuster, A.
+W.
+Senior, M.
+J.
+Shearn, A.
+Shorter, N.
+Shutty, V.
+Shvarts, S.
+Singh, V.
+Sivak, J.
+Skruzny, S.
+Small, V.
+Smelyanskiy, W.
+C.
+Smith, R.
+D.
+Somma, S.
+Springer, G.
+Sterling, D.
+Strain, J.
+Suchard, A.
+Szasz, A.
+Sztein, D.
+Thor, A.
+Torres, M.
+M.
+Torunbalci, A.
+Vaishnav, J.
+Vargas, S.
+Vdovichev, G.
+Vidal, B.
+Villalonga, C.
+V.
+Heidweiller, S.
+Waltman, S.
+X.
+Wang, B.
+Ware, K.
+Weber, T.
+White, K.
+Wong, B.
+W.
+K.
+Woo, C.
+Xing, Z.
+J.
+Yao, P.
+Yeh, B.
+Ying, J.
+Yoo, N.
+Yosri, G.
+Young, A.
+Zalcman, Y.
+Zhang, N.
+Zhu, and N.
+Zobrist, Quantum error correction below the surface code threshold (2024), arXiv:2408.13687 [quant-ph].
 - Koch et al.
@@
 Deng, Titanium Nitride Film on Sapphire Substrate with Low Dielectric Loss for Superconducting Qubits, Physical Review Applied 19, 024013 (2023).
+- Sivak et al.
+[2023] V.
+V.
+Sivak, A.
+Eickbusch, B.
+Royer, S.
+Singh, I.
+Tsioutsios, S.
+Ganjam, A.
+Miano, B.
+L.
+Brock, A.
+Z.
+Ding, L.
+Frunzio, S.
+M.
+Girvin, R.
+J.
+Schoelkopf, and M.
+H.
+Devoret, Real-time quantum error correction beyond break-even, Nature 616, 50 (2023).
 - Biznárová et al.
@@
 Kippenberg, Mechanically induced correlated errors on superconducting qubits with relaxation times exceeding 0.4 ms, Nature Communications 15, 3950 (2024).
+- Dunsworth et al.
+[2017] A.
+Dunsworth, A.
+Megrant, C.
+Quintana, Z.
+Chen, R.
+Barends, B.
+Burkett, B.
+Foxen, Y.
+Chen, B.
+Chiaro, A.
+Fowler, R.
+Graff, E.
+Jeffrey, J.
+Kelly, E.
+Lucero, J.
+Y.
+Mutus, M.
+Neeley, C.
+Neill, P.
+Roushan, D.
+Sank, A.
+Vainsencher, J.
+Wenner, T.
+C.
+White, and J.
+M.
+Martinis, Characterization and reduction of capacitive loss induced by sub-micron Josephson junction fabrication in superconducting qubits, Applied Physics Letters 111, 022601 (2017).
+- Wang et al. [2019] Z. Wang, S. Shankar, Z. Minev, P. Campagne-Ibarcq, A. Narla, and M. Devoret, Cavity Attenuators for Superconducting Qubits, Physical Review Applied 11, 014031 (2019).
+- Nersisyan et al.
+[2019] A.
+Nersisyan, E.
+A.
+Sete, S.
+Stanwyck, A.
+Bestwick, M.
+Reagor, S.
+Poletto, N.
+Alidoust, R.
+Manenti, R.
+Renzas, C.-V.
+Bui, K.
+Vu, T.
+Whyland, and Y.
+Mohan, Manufacturing low dissipation superconducting quantum processors, in 2019 IEEE International Electron Devices Meeting (IEDM) (IEEE, San Francisco, CA, USA, 2019) pp.
+31.1.1–31.1.4.
+- Wei et al.
+[2020] K.
+X.
+Wei, I.
+Lauer, S.
+Srinivasan, N.
+Sundaresan, D.
+T.
+McClure, D.
+Toyli, D.
+C.
+McKay, J.
+M.
+Gambetta, and S.
+Sheldon, Verifying multipartite entangled Greenberger-Horne-Zeilinger states via multiple quantum coherences, Physical Review A 101, 032343 (2020).
+- Moskalev et al.
+[2023] D.
+O.
+Moskalev, E.
+V.
+Zikiy, A.
+A.
+Pishchimova, D.
+A.
+Ezenkova, N.
+S.
+Smirnov, A.
+I.
+Ivanov, N.
+D.
+Korshakov, and I.
+A.
+Rodionov, Optimization of shadow evaporation and oxidation for reproducible quantum Josephson junction circuits, Scientific Reports 13, 4174 (2023).
+- Pishchimova et al.
+[2023] A.
+A.
+Pishchimova, N.
+S.
+Smirnov, D.
+A.
+Ezenkova, E.
+A.
+Krivko, E.
+V.
+Zikiy, D.
+O.
+Moskalev, A.
+I.
+Ivanov, N.
+D.
+Korshakov, and I.
+A.
+Rodionov, Improving Josephson junction reproducibility for superconducting quantum circuits: junction area fluctuation, Scientific Reports 13, 6772 (2023).
+- Pop et al.
+[2014] I.
+M.
+Pop, K.
+Geerlings, G.
+Catelani, R.
+J.
+Schoelkopf, L.
+I.
+Glazman, and M.
+H.
+Devoret, Coherent suppression of electromagnetic dissipation due to superconducting quasiparticles, Nature 508, 369 (2014).
 - Gyenis et al.
@@
 Sunada, H.
-Kivijärvi, L.
+Kivijärvi, J.
+Albanese, L.
 Grönberg, J.-P.
@@
 Govenius, and M.
-Möttönen, Data and GDS file for ”Methods to achieve near-millisecond energy relaxation and dephasing times for a superconducting transmon qubit (2024).
+Möttönen, Data and GDS file for ”Methods to achieve near-millisecond energy relaxation and dephasing times for a superconducting transmon qubit”missing (2024).
 - Burnett et al.
@@
 Möttönen, Unimon qubit, Nature Communications 13, 6895 (2022).
-Appendix A Extended Data Figures
+Liite A Extended Data Figures and Tables
+| Ref. | (GHz) | (µs) | (106) | (µs) |
+| 9 | 3.73 | 303 | 7.1 | 201 |
+| 10 | 3.92 | 401 | 9.9 | N/A |
+| 11 | 3.86 | 393 | 9.5 | N/A |
+| 12 | 4.24 | 292 | 7.7 | N/A |
+| 13 | 5.92 | 280 | 10.4 | 238 |
+| 14 | 3.02 | 270 | 5.1 | 307 |
+| 15 | 3.95 | 323 | 8.0 | 286 |
+| 16 | 4.79 | 210 | 6.5 | 290 |
+| This work | 2.89 | 502 | 9.1 | 541 |
+| Equipment | Description |
+| Sputtering system | Eclipse Mark IV |
+| Maskless aligner | Heidelberg Instruments MLA150 |
+| Plasma processing system | Oxford Instruments Plasmalab |
+| 80 Plus | |
+| Electron-beam writer | Raith EBPG5200 |
+| Electron-beam evaporator | Plassys MEB700S2-III |
+| Vapor prime oven | YES III |
+| Equipment | Description | Abbreviation |
+| Dilution refrigerator | Bluefors XLD400sl | |
+| RFSoC evaluation board | Xilinx Zynq UltraScale+ RFSoC ZCU216 with add-ons CLK104 and XM655 | |
+| Frequency standard | Stanford Research Systems FS725 10-MHz Rb Standard | |
+| Coaxial cables | Totoku TCF258AA2000/1500 (room temperature) | |
+| Mini-Circuits 141-72SM+ (room temperature) | ||
+| Bluefors 0.86 mm SCuNi-CuNi semi-rigid (cryogenic input lines) | ||
+| Bluefors 0.86 mm NbTi-NbTi semi-rigid (cryogenic output lines, MXC to 4 K) | ||
+| Bluefors 2.19 mm SCuNi-CuNi semi-rigid (cryogenic output lines, 4 K to RT) | ||
+| HEMT amplifiers | Low Noise Factory LNF-LNR4_8F (room temperature) | HEMT |
+| Low Noise Factory LNF-LNC4_8F (cryogenic) | HEMT∗ | |
+| Low pass filters | RLC F-30-8000-R (3-dB point = 8.4 GHz, 60-dB stopband = 10.8–40 GHz) | LPF |
+| Mini-Circuits VLF-5850+ (3-dB point = 6.54 GHz) | LPF∗ | |
+[…]
+| Bandpass filters | Mini-Circuits VBFZ-3590+ (2-dB passband = 3–4.3 GHz) | VBFZ |
+| Mini-Circuits VBFZ-6260-S+ (2-dB passband = 5.6–7 GHz) | VBFZ∗ | |
+| Highpass filter | Mini-Circuits VHF-8400+ (20-dB stopband = DC–7.8 GHz) | HPF |
+| Eccosorb filter | Kawashima Manufacturing 3.5mm Filter Connector L9.6 (0.32 dB/GHz) | Ecco. |
+| Attenuator | Bluefors cryogenic attenuators | dB |
+| XMA 2082-6418-20-CRYO | 20 dB∗ | |
+| Mini-Circuits BW-S10-2W263+ (room temperature) | 10 dB† | |
+| Isolators | Low Noise Factory LNF-ISISC4_8A (4–8 GHz) | |
+| Low Noise Factory LNF-ISISC4_12A (4–12 GHz) | ||
+| Circulator | Low Noise Factory LNF-CIC4_12A (4–12 GHz) | |
+| TWPA | Supplied by VTT (Technical Research Center of Finland Ltd.) | TWPA |
+| TWPA pump generator | Valon 5015 Frequency Synthesizer | |
+| Voltage source | Stanford Research Systems SIM928 Isolated Voltage Source | V |
+| Sample holder | QDevil QCage.24 with QCage Magnetic Shielding | QCage.24 |